# SimROD: A Simple Baseline for Raw Object Detection with Global and Local Enhancements

### v2

@@ 4.3 Comparison with State-of-the-Art Methods
Table 3: Segmentation Performance on ADE20K-Raw [45, 8]. Comparison of semantic segmentation (mIoU) results under normal, overexposure (Over-exp), and low-light (Dark) conditions using Segformer [37]. Bold numbers indicate the best results for each condition.
@@
Demosacing [8] 47.47 45.69 37.55 0.00
Karaimer et Brown [8, 18] 45.48 42.85 37.32 0.00
Lite-ISP [8, 43] 43.22 42.01 5.52 0.00
InvISP [8, 38] 47.82 44.30 4.03 0.00
SID [8, 6] - - 37.06 0.00
DNF [8, 17] - - 35.88 0.00
Dirty-Pixel [8, 9] 47.86 46.50 38.02 4.28
RAW-Adapter [8] 47.95 46.62 38.75 0.30
@@
Table 4: Ablation Study on the Impact of Different Components over Pascal-Raw [27] and LOD [15] datasets with YoloX-Tiny [44]. The term ‘Guided Channel’ refers to the use of a specific channel for local enhancement in GGLE, while GGLE without a guided channel indicates a configuration with only the RGGB branch. Bold numbers indicate the best performance.
Pretrained GGE GGLE Guided LOD [15] Pascal-Raw [27]
Weight Channel AP50 AP50
@@
GG 46.3 95.1
Table 5: Comparison of DIAP [39] and our GGE on ROD [39], Pascal-Raw [27], and LOD [15] datasets with YoloX-Tiny [44]. Results are shown in AP and AP50 for each dataset, with additional parameters and GFLOPs also provided.
Dataset Metric DIAP [39] Our GGE
ROD [39] AP 30.7 30.9
@@
4.4 Analysis and Discussion
All experiments in this section follow the settings of DIAP [39], unless stated otherwise.
@@ 4.4 Analysis and Discussion
Table 4 presents an ablation study on the LOD [15] and Pascal-Raw [27] datasets using YoloX-Tiny [44] to evaluate the impact of different enhancement modules. The best AP and AP50 results are achieved when all four components are used, indicating the combined enhancements’ positive effect on model performance. Moreover, the green channel guidance outperforms the red and blue guidance, indicating the unique value of the green channel. Notably, using the red and blue channel guidance even leads to the performance drop on LOD [15] which is low-light, noisy, and more challenging. We also evaluate RGGB Table 4. We find that that solely using the Green Channels (GG) alone outperforms the combination of all channels (RGGB). This is attributed to the significant noise in R and B of RGGB, which affects the learning of the model. The inferior performance of RB compared to R alone further corroborates this, suggesting that combining all channels as guidance does not yield better results.
Table 6: Effect of Green Channel Sampling Frequency. Reducing the green channel sampling frequency to 0.5$\times$ leads to a significant performance drop, especially on the low-light LOD dataset, highlighting the importance of dense green channel information.
G Sampling Frequency LOD [15] Pascal-Raw [27, 8]
@@
0.5$\times$ (reduced) 22.0 40.3 67.5 94.3
1$\times$ (default) 24.3 42.0 67.9 94.2
RAW Images w. GT
RAW Distribution
DIAP [39] Det.
Distribution after DIAP [39]
Our SimROD Det.
Distribution after Our SimROD
Figure 4: Visualization on the LOD [15] Dataset. We present detection and distribution results processed with DIAP [39] and our SimROD. The figure includes RAW images with detection annotations, detection results, and RGB pixel distributions for both methods. The pixel distribution of the enhanced images from our SimROD more closely approximates a normal distribution, which facilitates more efficient feature learning for neural networks.
GGE v.s. DIAP [39]
We compares DIAP [39] and our GGE across ROD [39], Pascal-Raw [27], and LOD [15] datasets in Table 5. Our GGE achieves comparable or slightly improved performance in certain metrics while reducing parameters and GFLOPs, reflecting a more efficient model design. Note that GGE only contains four learnable parameters.
@@
We investigate how the green channel’s sampling frequency affects RAW object detection performance (Table 6). In our experiments, we feed the green channel into DIAP [39] and compare two strategies: the default 1× sampling based on the RGGB pattern versus a reduced 0.5× sampling (using only one green value). The results show that lowering the green channel frequency leads to a significant performance drop, especially on the low-light LOD dataset. This confirms that the denser spatial information provided by the green channel—due to its higher frequency in the Bayer pattern—is crucial for robust performance.
4.5 Visualization
Figure 4 provides a comprehensive visualization of detection results and channel distributions for both the DIAP [39] and our SimROD across various scenes. A clear pattern emerges from these visualizations: our method consistently produces more accurate and stable detection results compared to DIAP [39]. Notably, the pixel distribution of the enhanced images more closely approximates a normal distribution. This characteristic facilitates more efficient feature learning for neural networks, reducing the impact of outliers and thus improving detection accuracy. More visualizations are provided in the supplementary material.
5 Related Work
Object Detection
Object detection has been an active area of research in computer vision [10, 20, 36]. Early works[11] relied on manually crafted features to detect objects. With the advent of deep learning, convolutional neural networks (CNNs) revolutionized object detection [13, 12, 32]. They use CNNs to extract features from proposed regions, significantly improving accuracy. Single-stage detectors, such as Yolo family [31, 30, 34, 35] and SSD [22], aimed to reduce inference time by directly predicting bounding boxes and class probabilities from the entire image without relying on region proposals. Recently, transformer-based approaches [46, 5] have demonstrated that object detection can benefit from attention mechanisms.
@@ 5 Related Work
RAW data is the input for the image signal processor (ISP)[39, 18, 26, 29]. Considerable efforts have been made to utilize the RAW data to improve detection performance and robustness. Hong et al. [15] designed an auxiliary task for image reconstruction on synthetic datasets to improve detection performance. Onzon et al. [28] proposed estimating camera exposure parameters from previous frames to achieve well-exposed images. Morawski et al. [23] designed an ISP pipeline called GenISP, which applies learnable white balance and color correction transformations to RAW data. Yoshimura et al. [40] decomposed control into the entire dataset, along with fine-tuning for individual images. They introduced a latent update style controller to manage the stages of a differentiable ISP. Xu et al. [39] developed an image-level adjustment module and a pixel-level adjustment module to learn transformations. Chen et al. [7] proposed an activation function to extract features from RAW data effectively. Cui et al. [8] combined an input-level adapter with a model-level adapter to enhance downstream task performance. Additionally, Yoshimura et al. [41] proposed using noise-accounted raw augmentation to enhance recognition performance. Despite these advancements, most of these approaches often rely on simulating expert-tuning ISP stages, resulting in redundant parameters and increased computational complexity, which can adversely affect both accuracy and efficiency.
@@
Many works have focused on enhancing images from RAW data. Brooks et al. [3] introduced a method to transform sRGB images into synthetic RAW data and trained a neural network model for denoising. Zou et al. [47] suggested separating challenging and easier regions and using dual intensity and global spatial guidance to reconstruct images from RAW images. Guo et al. [14] proposed leveraging the green channel as a prior to jointly perform demosaicking and denoising on images, demonstrating the unique value of the green channel. However, these methods depend on paired datasets, which, even when artificially synthesized, often result in suboptimal data that may not be directly applicable to object detection tasks.
6 Conclusion
Supplementary Material
In this supplementary material, we present:
  • Section A: Segmentation Performance on the Real-World iPhone XSmax [19] Dataset.
  • Section B: More analysis of our Global Gamma Enhancement(GGE), including a brief introduction to the Gamma Transformation, hyperparameters sensitivity analysis, ablation study of GGE and comparision between sRGB and RAW.
  • Section C. We compared the inference time with DIAP [39] and RAW-Adapter [8], further demonstrating the efficiency of our approach.
  • Section D: More visualization, including additional comparison of our SimROD and DIAP [39] on the detection dataset, comparison of SimROD and RAW-Adapter [8] on the detection dataset, and comparison of SimROD and RAW-Adapter [8] on the segmentation dataset.
A Segmentation on Real-World iPhone XSmax Dataset
The iPhone XSmax [19] is a real-world RAW semantic segmentation dataset that consists of 1153 RAW images with their corresponding semantic labels, where 806 images are set as training set and the other 347 images are set as evaluation set. Following Raw-adapter [8], we adopt Segformer[37] framework with MIT-B5 backbone, training iterations are set to 20000, and other settings are same as ADE 20K RAW [8]’s setting, as defined by  [8]. As shown in Table 7, the SimROD could also achieve superior results with its parameter efficiency.
Table 7: Segmentation Performance on the Real-World iPhone XSmax [19] Dataset. Bold number indicates the best result.
@@ A Segmentation on Real-World iPhone XSmax Dataset
Karaimer et Brown [18] 60.41 0.00
LiteISP [43] 59.84 0.00
Dirty-Pixel [9] 61.85 4.28
RAW-Adapter [8] 62.45 0.30
@@
B More analysis of our GGE
B.1 Gamma Transformation
@@
B.2 Hyperparameters Sensitivity Analysis
The hyperparameters $\gamma_{min}$ and $\gamma_{max}$ are defined in our Global Gamma Enhancement (GGE). We have set their values by default according to  [39]. To further analyze the GGE module, we have conducted hyperparameter sensitivity experiments on both LOD [15] and Pascal-Raw [27].
@@ B.2 Hyperparameters Sensitivity Analysis
(a) Effect of Gamma Range on LOD [15].
(b) Effect of Gamma Range on Pascal-Raw [27].
B.3 Ablation Study
The results presented in Table 8 emphasize the importance of Gamma transformation techniques for object detection on the ROD [39] dataset, which features significant pixel intensity variations between nighttime and daytime scenes. Our GGE substantially improves the mAP and $\text{AP}_{50}$ scores compared to settings without GGE, indicating its effectiveness in reducing the variance of pixel intensity and alleviating challenges arising from drastic changes in lighting conditions. When GGE is combined with our Green-Guided Local Enhancement (GGLE), the model achieves the highest performance. This demonstrates that Gamma transformations are essential for achieving robust detection performance in datasets with extreme lighting variability.
Table 8: Ablation Study on ROD [39]. The proposed GGE significantly boosts detection performance, particularly when combined with the proposed GGLE. Bold numbers indicate the best result.
GGE GGLE mAP $\text{AP}_{50}$
4.8 9.5
30.9 53.5
1.3 3.6
33.1 57.6
B.4 Comparison between sRGB and RAW.
@@ B.3 Ablation Study
Table 9: Comparison between sRGB and RAW on LOD[15] and Pascal-Raw [27].
Data GGE GGLE LOD[15] Pascal-Raw [27]
@@
C Latency
As shown in Tab. 10, we compare our method with DIAP [39] and Raw-Adaptor [8] on mAP on the LOD [15] dataset, parameter count, and latency. Our method achieves higher mAP while drastically reducing parameters and inference latency across different architectures. These results demonstrate the clear advantage of our approach in efficiency and performance.
Table 10: Comparison of Latency and Parameters on the LOD [15].
@@ C Latency
D More Visualization Results
Comparison of our SimROD and DIAP [39] on the detection dataset, comparison of SimROD and RAW-Adapter [8] on the detection dataset, and comparison of SimROD and RAW-Adapter [8] on the segmentation dataset are provided in Figure 6, Figure 7 and Figure 8 respectively.
@@ D More Visualization Results
DIAP [39] Det.
Distribution after DIAP [39]
@@
Figure 6: Visualization on the Object Detection Dataset. We present detection and distribution results processed with DIAP [39] and our SimROD. The figure includes RAW images with detection annotations, detection results, and RGB pixel distributions for both methods. The pixel distribution of the enhanced images from our SimROD more closely approximates a normal distribution, which facilitates more efficient feature learning for neural networks.
RAW Images w. GT
RAW Distribution
RAW-Adapter [8] Det.
Distribution after RAW-Adapter [8]
Our SimROD Det.
Distribution after Our SimROD
Figure 7: Visualization on the Object Detection Dataset. We present detection and distribution results processed with RAW-Adapter [8] and our SimROD. The figure includes RAW images with detection annotations, detection results, and RGB pixel distributions for both methods.
@@
Figure 8: Semantic segmentation visualization results on ADE 20K RAW [8].

### v3

@@ Discussion
Table 3: Segmentation Performance on ADE20K-Raw (Zhou et al. 2017; Cui and Harada 2024). Comparison of semantic segmentation (mIoU) results under normal, overexposure (Over-exp), and low-light (Dark) conditions using Segformer (Xie et al. 2021). Bold numbers indicate the best results for each condition.
@@
Demosacing 47.47 45.69 37.55 0.00
Karaimer et Brown 45.48 42.85 37.32 0.00
Lite-ISP 43.22 42.01 5.52 0.00
InvISP 47.82 44.30 4.03 0.00
SID - - 37.06 0.00
DNF - - 35.88 0.00
Dirty-Pixel 47.86 46.50 38.02 4.28
RAW-Adapter 47.95 46.62 38.75 0.30
@@
Table 4: Ablation Study on the Impact of Different Components over Pascal-Raw (Omid-Zohoor, Ta, and Murmann 2014) and LOD (Hong et al. 2021) datasets with YoloX-Tiny (Zheng et al. 2021). The term ‘Guided Channel’ refers to the use of a specific channel for local enhancement in GGLE, while GGLE without a guided channel indicates a configuration with only the RGGB branch. Bold numbers indicate the best performance.
Pretrained GGE GGLE GGLE LOD Pascal-Raw
Weight (RGGB) (Guidance) AP50 AP50
@@
$G$ 46.3 95.1
Table 5: Comparison of DIAP (Xu et al. 2023) and our GGE on ROD (Xu et al. 2023), Pascal-Raw (Omid-Zohoor, Ta, and Murmann 2014), and LOD (Hong et al. 2021) datasets with YoloX-Tiny (Zheng et al. 2021). Results are shown in AP and AP50 for each dataset, with additional parameters and GFLOPs also provided.
Dataset Metric DIAP Our GGE
ROD AP 30.7 30.9
@@
Analysis and Discussion
@@
Table 4 presents an ablation study on the LOD (Hong et al. 2021) and Pascal-Raw (Omid-Zohoor, Ta, and Murmann 2014) datasets using YoloX-Tiny (Zheng et al. 2021) to evaluate the impact of different enhancement modules. The best AP and AP50 results are achieved when all four components are used, indicating the combined enhancements’ positive effect on model performance. Moreover, the green channel guidance outperforms the red and blue guidance, indicating the unique value of the green channel. Notably, using the red and blue channel guidance even leads to the performance drop on LOD (Hong et al. 2021) which is low-light, noisy, and more challenging. We also evaluate RGGB Table 4. We find that that solely using the Green Channels (GG) alone outperforms the combination of all channels (RGGB). This is attributed to the significant noise in R and B of RGGB, which affects the learning of the model. The inferior performance of RB compared to R alone further corroborates this, suggesting that combining all channels as guidance does not yield better results.
[…]
G Sampling Frequency LOD Pascal-Raw
@@
[…]
GGE v.s. DIAP (Xu et al. 2023)
We compares DIAP (Xu et al. 2023) and our GGE across ROD (Xu et al. 2023), Pascal-Raw (Omid-Zohoor, Ta, and Murmann 2014), and LOD (Hong et al. 2021) datasets in Table 5. Our GGE achieves comparable or slightly improved performance in certain metrics while reducing parameters and GFLOPs, reflecting a more efficient model design. Note that GGE only contains four learnable parameters.
@@
We investigate how the green channel’s sampling frequency affects RAW object detection performance (Table 6). In our experiments, we feed the green channel into DIAP (Xu et al. 2023) and compare two strategies: the default 1× sampling based on the RGGB pattern versus a reduced 0.5× sampling (using only one green value). The results show that lowering the green channel frequency leads to a significant performance drop, especially on the low-light LOD dataset. This confirms that the denser spatial information provided by the green channel—due to its higher frequency in the Bayer pattern—is crucial for robust performance.
The Value of $\gamma$ in Training .
As shown in Fig. 4, the $\gamma$ values slightly increase during training leading to the resulting pixel values are slightly reduced. This is reasonable, as the initial value of 0.114 is lower than the commonly used default of 1/2.2 in standard gamma correction.
Figure 4: $\gamma$ across epochs on LOD and Pascal-Raw.
Conclusion
Related Work
@@ Conclusion
Object detection has been an active area of research in computer vision (Everingham et al. 2010; Lin et al. 2014; Wu, Sahoo, and Hoi 2020; Tian et al. 2025; Ma et al. 2024; Wang et al. 2025b, a). RAW data serves as the input to the image signal processor (ISP) and has attracted significant attention due to its unique value.(Xu et al. 2023; Karaimer and Brown 2016; Nishimura et al. 2018; Ramanath et al. 2005). Considerable efforts have been made to utilize the RAW data to improve detection performance and robustness. (Hong et al. 2021) designed an auxiliary task for image reconstruction on synthetic datasets to improve detection performance. (Onzon, Mannan, and Heide 2021) proposed estimating camera exposure parameters from previous frames to achieve well-exposed images. (Morawski et al. 2022) designed an ISP pipeline called GenISP, which applies learnable white balance and color correction transformations to RAW data. (Yoshimura et al. 2023a) decomposed control into the entire dataset, along with fine-tuning for individual images. They introduced a latent update style controller to manage the stages of a differentiable ISP. (Xu et al. 2023) developed an image-level adjustment module and a pixel-level adjustment module to learn transformations. (Chen, Tai, and Ma 2024) proposed an activation function to extract features from RAW data effectively. (Cui and Harada 2024) combined an input-level adapter with a model-level adapter to enhance downstream task performance. Additionally, (Yoshimura et al. 2023b) proposed using noise-accounted raw augmentation to enhance recognition performance. Despite these advancements, most of these approaches often rely on simulating expert-tuning ISP stages, resulting in redundant parameters and increased computational complexity, which can adversely affect both accuracy and efficiency.
@@
Many works have focused on enhancing images from RAW data. (Brooks et al. 2019) introduced a method to transform sRGB images into synthetic RAW data and trained a neural network model for denoising. (Zou, Yan, and Fu 2023) suggested separating challenging and easier regions and using dual intensity and global spatial guidance to reconstruct images from RAW images. (Guo, Liang, and Zhang 2021) proposed leveraging the green channel as a prior to jointly perform demosaicking and denoising on images, demonstrating the unique value of the green channel. However, these methods depend on paired datasets, which, even when artificially synthesized, often result in suboptimal data that may not be directly applicable to object detection tasks.
Acknowledgement
We thank Caizhi Zhu, Yinqiang Zheng, Xuanlong Yu, Zechao Hu, Hao Li, Zhengwei Yang, Song Ouyang, Jingyu Xu, Likai Tian, and Runqi Wang for their valuable support. This work was funded by the National Natural Science Foundation of China (Grant No. 62571379). The numerical calculations in this paper have been done on the supercomputing system in the Supercomputing Center of Wuhan University.
Appendix
In this appendix, we present:
  • Segmentation Performance on the Real-World iPhone XSmax (Li et al. 2024) Dataset.
  • Detection Performance on Synthetic Over-exposure and Dark Datasets.
  • Visualization, including additional comparison of our SimROD and DIAP (Xu et al. 2023) on the detection dataset, comparison of SimROD and RAW-Adapter (Cui and Harada 2024) on the detection dataset, and comparison of SimROD and RAW-Adapter (Cui and Harada 2024) on the segmentation dataset.
  • More analysis of our Global Gamma Enhancement(GGE), including a brief introduction to the Gamma Transformation, hyperparameters sensitivity analysis, ablation study of GGE and comparision between sRGB and RAW.
  • We compared the inference time with DIAP (Xu et al. 2023) and RAW-Adapter (Cui and Harada 2024), further demonstrating the efficiency of our approach.
(a) Effect of Gamma Range on LOD (Hong et al. 2021).
(b) Effect of Gamma Range on Pascal-Raw (Omid-Zohoor, Ta, and Murmann 2014).
Segmentation on Real-World iPhone XSmax Dataset
The iPhone XSmax (Li et al. 2024) is a real-world RAW semantic segmentation dataset that consists of 1153 RAW images with their corresponding semantic labels, where 806 images are set as training set and the other 347 images are set as evaluation set. Following Raw-adapter (Cui and Harada 2024), we adopt Segformer(Xie et al. 2021) framework with MIT-B5 backbone, training iterations are set to 20000, and other settings are same as ADE 20K RAW (Cui and Harada 2024)’s setting, as defined by  (Cui and Harada 2024). As shown in Table 7, the SimROD could also achieve superior results with its parameter efficiency.
Table 7: Segmentation Performance on the Real-World iPhone XSmax (Li et al. 2024) Dataset. Bold number indicates the best result.
@@ Acknowledgement
Karaimer et Brown 60.41 0.00
LiteISP 59.84 0.00
Dirty-Pixel 61.85 4.28
RAW-Adapter 62.45 0.30
@@
Detection Performance on Synthetic Over-exposure and Dark Datasets
Following  (Cui and Harada 2024), we also report results on the Over-exp and Dark versions in Tab. 8, where SimROD continues to outperform existing methods by a large margin.
Table 8: Over-exposure and dark version Pascal-Raw.
Method Over-exposure Dark
Dirty-Pixel [8,9] 88.0 80.8
RAW-Adapter 88.7 82.5
SimROD 89.4 84.1
Visualization
Figure 6 provides a comprehensive visualization of detection results and channel distributions for both the DIAP (Xu et al. 2023) and our SimROD across various scenes. A clear pattern emerges from these visualizations: our method consistently produces more accurate and stable detection results compared to DIAP (Xu et al. 2023). Notably, the pixel distribution of the enhanced images more closely approximates a normal distribution. This characteristic facilitates more efficient feature learning for neural networks, reducing the impact of outliers and thus improving detection accuracy.
Comparison of SimROD and RAW-Adapter on the segmentation dataset are provided in Figure 7.
More analysis of our GGE
Gamma Transformation
@@
Hyperparameters Sensitivity Analysis
The hyperparameters $\gamma_{min}$ and $\gamma_{max}$ are defined in our Global Gamma Enhancement (GGE). We have set their values by default according to  (Xu et al. 2023). To further analyze the GGE module, we have conducted hyperparameter sensitivity experiments on both LOD (Hong et al. 2021) and Pascal-Raw (Omid-Zohoor, Ta, and Murmann 2014).
@@
Ablation Study
The results presented in Table 9 show that both proposed components—GGE and GGLE—exhibit consistent behavior across all evaluated input resolutions ($\times640$, $\times960$, and $\times1280$). Crucially, using GGLE alone leads to near-complete failure (e.g., mAP $\approx$ 0), indicating that local enhancement without global intensity normalization is ineffective. In contrast, GGE alone already provides substantial improvements over the baseline, and further combining it with GGLE consistently yields the best performance. This demonstrates that GGE is essential for object detection on the ROD (Xu et al. 2023) dataset, while GGLE is only beneficial when built upon GGE. To better utilize GPU memory capacity, the experiments reported in this table were conducted using 4 GPUs with a total batch size of 16 (i.e., 4 per GPU), which slightly differs from the hyperparameters used in the main text. The necessity of GGE can be explained by the structure of ROD: daytime and nighttime images are interleaved in the same dataset, resulting in extreme pixel intensity discrepancies. Without a global adaptive normalization like GGE, the model cannot reconcile these divergent distributions, rendering subsequent local enhancements (e.g., GGLE) ineffective or even harmful.
Table 9: Ablation Study on ROD (Xu et al. 2023). Bold numbers indicate the best result.
Input Size GGE GGLE mAP AP50
$\times1280$ 49.5 76.7
0.0 0.0
51.5 78.9
$\times960$ 43.8 70.9
0.0 0.2
45.6 73.0
$\times640$ 32.4 56.3
0.1 0.4
34.8 59.5
Comparison between sRGB and RAW.
@@
Table 10: Comparison between sRGB and RAW on LOD(Hong et al. 2021) and Pascal-Raw (Omid-Zohoor, Ta, and Murmann 2014).
Data GGE GGLE LOD Pascal-Raw
@@
Latency
As shown in Tab. 11, we compare our method with DIAP (Xu et al. 2023) and Raw-Adaptor (Cui and Harada 2024) on mAP on the LOD (Hong et al. 2021) dataset, parameter count, and latency. Our method achieves higher mAP while drastically reducing parameters and inference latency across different architectures. These results demonstrate the clear advantage of our approach in efficiency and performance.
Table 11: Comparison of Latency and Parameters on the LOD.
@@
DIAP (Xu et al. 2023) Det.
Distribution after DIAP (Xu et al. 2023)
@@
Figure 6: Visualization on the Object Detection Dataset. We present detection and distribution results processed with DIAP (Xu et al. 2023) and our SimROD. The figure includes RAW images with detection annotations, detection results, and RGB pixel distributions for both methods. The pixel distribution of the enhanced images from our SimROD more closely approximates a normal distribution, which facilitates more efficient feature learning for neural networks.
@@
Figure 7: Semantic segmentation visualization results on ADE 20K RAW (Cui and Harada 2024).